# Synthesizing Programmatic Reinforcement Learning Policies with Large Language Model Guided Search

### v2

@@ Appendix B Karel DSL details
There are 5 symbols in the Karel DSL: statement, condition, action, boolean, and number. The production rules and Python converting rules are listed in Table 4. The statement counts the mean total usage of converting one statement into two statements at the root of the abstract syntax tree (AST), and the statement counts are limited to 6 in search of optimal programs. The program length is the total number of tokens used in string representation and is set to 44. The depth in DSL is the recursive call of control flows, which is limited to 4. These limitations are the same as the original settings in LEAPS [Trivedi et al., 2021]. The probabilities of the production rules are listed in Table 5. The mutation process selects one random node in the AST. We use the production rules to sample a sub-tree to replace the original node. The Karel syntax has many symbol-related parentheses, e.g., “i(” and “i)” which are related to the “IF” and “IFELSE”, thus LLM sometimes makes mistakes translating Python to DSL. We list common mistakes LLM made in Table 6 and how we solve these problems.
@@ Appendix C Karel tasks
We give the large language model 5 pieces of information related to the tasks: task name, map description, initial Position, task goal, and task return. By converting the original task description into these 5 categories, this design allows the user to easily fill in all categories for new tasks. Here we list all the information pieces for these tasks. Section I.1.2 shows how these 5 different components are filled into the placeholders in the user prompt. Section D.2 show all filled results with the task DoorKey. There are three sets of tasks: Karel [Trivedi et al., 2021], Karel-Hard [Liu et al., 2023], and new Karel tasks. Section C.1 lists all 6 Karel tasks, Section C.2 lists all 4 Karel-Hard tasks, and Section C.3 lists the 2 new tasks. Example figures of Karel tasks are in Figure 10, Karel-Hard tasks are in Figure 11 and new Karel tasks are in Figure 12.
@@ Appendix F Baselines and their hyperparameters
We compare our framework to four baselines, LEAPS [Trivedi et al., 2021], HPRL [Liu et al., 2023], CEBS [Carvalho et al., 2024], and HC [Carvalho et al., 2024]. LEAPS, HPRL, and CEBS search programs in the latent spaces. HC search program in the programmatic space. The pseudo-code of CEBS is described in Algorithm 1, and the pseudo-code of HC is described in Algorithm 2.
F.1 LEAPS
@@
F.2 Hierarchical programmatic reinforcement learning (HPRL)
@@
F.3 Cross-entropy beam search (CEBS)
CEBS [Carvalho et al., 2024] also utilizes the pre-trained VAE weight from LEAPS [Trivedi et al., 2021] to search program in the latent space. All the hyperparameter follows the original CEBS with neighborhood $K$ equal to 64, candidate $E$ equal to 16, and noise ratio $\sigma$ equal to 0.25 for all ten Karel tasks. The pseudo-code is described in Algorithm 1.
F.4 Hill climbing (HC)
@@ Appendix H Aggregated performance across all tasks
Appendix I LLM prompts for ablations and revision
@@
Ablation prompts for program generation methods. We conduct an ablation study in Section 5.3 to justify that our LLM-generating program method performs best in acceptance rate and best return. Here we list the complete prompts. There are three approaches to generating DSL programs: Pythonic-DSL, Python, and DSL. There are two types of prompt, system prompt and user prompt, in all approaches. In the system prompt, both the Python and the Pythonic-DSL approaches contain limitations of Python usage. The DSL approach contains the grammar of the Karel, and the Pythonic-DSL contains the paired Python-like and Karel production rules. All of the system prompts contain the environment physics, action, and perception. On the other hand, the user prompt contains five placeholders for task name, map description, initial position, task goal, and task reward. In Appendix C, all of the task-dependent information can be filled in the placeholders. Section D.2 shows the full Pythonic-DSL prompt with the task DoorKey filled in the placeholders. The Pythonic-DSL prompts are in Section I.1, and the Python prompts are in Section I.2, and the DSL prompts are in Section I.3.
@@
Prompts for LLM revision. We list all of the user prompts in the experiment of LLM revision, the system prompt is the same as the one in Section I.1. We implement four approaches to revising the program: Regenerate, Regenerate with reward, Agent execution trace, and Agent and program execution trace. We ask the LLM to regenerate with all generated programs in the last round without repetition, and the Regenerate prompt is in Section I.4. We ask the LLM to regenerate with the program relating to reward, and the revision prompt of the Regenerate with reward is in Section I.5. We utilize the execution trace of the Karel agent in the grid world, and the revision prompt of the Agent execution trace is in Section I.6. We provided the action/perception call and executing line, and the revision prompt of the Agent and program execution trace is in Section I.7. The program Figure 16 is the revision target of our method revision method in Section I.6 and Section I.7. Figure 16 is the trajectories of the program we used in method Section I.6 and Section I.7. In the example, the origin program has an average return of 0.5. LLM revision with agent trace can reach 0.640625, while LLM revision with both agent and program traces can reach a result of 0.8125.
@@
[!ht] {minted}python def run(): while noMarkersPresent(): if frontIsClear(): move() else: turnLeft() pickMarker() while frontIsClear(): move() turnRight() while noMarkersPresent(): if frontIsClear(): move() else: turnLeft() if frontIsClear(): move() turnRight() putMarker() The revision results of Section I.6.
@@
[!ht] {minted}python def run(): while noMarkersPresent(): if frontIsClear(): move() else: turnLeft() pickMarker() while noMarkersPresent(): if rightIsClear(): turnRight() move() else: if frontIsClear(): move() else: turnLeft() putMarker() The revision results of Section I.7.
I.1 Pythonic-DSL
I.1.1 System prompt
@@ I.1.2 User prompt
I.2 Python
I.2.1 System prompt
@@ I.4 Regenerate
I.5 Regenerate with reward
@@
I.6 Agent execution trace
@@
I.7 Agent and Program execution trace.
@@
Appendix J Limitation

### v3

@@ Appendix B Karel DSL details
In our framework, we ask the LLM to write the program in Python and then translate it into DSL, all in one response. We include a sample output from the LLM in Section D.3. To translate the responded program to an abstract syntax tree (AST), we use an off-the-shelf Karel parser which can convert a string in DSL into an AST. We implement a translator to convert a Python program into Karel DSL. We list common mistakes LLM made in Table 6 and how we solve these problems.
There are 5 symbols in the Karel DSL: statement, condition, action, boolean, and number. The production rules and Python converting rules are listed in Table 4. The statement counts are the total statements connected to the root node of the abstract syntax tree (AST), and the statement counts are limited to 6 in search of optimal programs. The program length is the total number of tokens used in string representation and is set to 44. The depth in DSL is the recursive call of control flows, which is limited to 4. These limitations are the same as the original settings in LEAPS [Trivedi et al., 2021]. The probabilities of the production rules are listed in Table 5. The mutation process selects one random node in the AST. We use the production rules to sample a sub-tree to replace the original node. The Karel syntax has many symbol-related parentheses, e.g., “i(” and “i)” which are related to the “IF” and “IFELSE”, thus LLM sometimes makes mistakes translating Python to DSL.
@@ Appendix C Karel tasks
We give the large language model 5 pieces of information related to the tasks: task name, map description, initial Position, task goal, and task return. By converting the original task description into these 5 categories, this design allows the user to easily fill in all categories for new tasks. Here we list all the information pieces for these tasks. Section K.1.2 shows how these 5 different components are filled into the placeholders in the user prompt. Section D.2 show all filled results with the task DoorKey. There are three sets of tasks: Karel [Trivedi et al., 2021], Karel-Hard [Liu et al., 2023], and new Karel tasks. Section C.1 lists all 6 Karel tasks, Section C.2 lists all 4 Karel-Hard tasks, and Section C.3 lists the 2 new tasks. Example figures of Karel tasks are in Figure 11, Karel-Hard tasks are in Figure 12 and new Karel tasks are in Figure 13.
@@ Appendix E Scheduled Hill Climbing detail
E.1 Design intuition
Scheduled HC comprises logarithmic interpolation, sinusoidal schedule, and logarithmic ratio. Our intuition to logarithmic interpolation and logarithmic ratio aims to allocate the most appropriate neighborhood size, $k$, to tasks of varying difficulties. The logarithm of the number of executed programs provides an indication of task difficulty, with the intuition that the optimal $k$ should increase exponentially according to the structure of AST. For the sinusoidal schedule, we prioritize maintaining a stable neighborhood size $k$ during both the early and final stages of the training process.
E.2 Variants of Scheduled Hill Climbing components
We add their linear counterpart, resulting in 2 variants for each component and 8 variants (2x2x2=8) on the whole.
E.2.1 Interpolation variants
  • Logarithmic interpolation
    $log_{2}{k(n)}$ $=(1-r(n))\log_{2}{K_{start}}+r(n)\log_{2}{K_{end}}$ (4)
  • Linear interpolation
    $k(n)$ $=(1-r(n))K_{start}+r(n)K_{end}$ (5)
E.2.2 Schedule variants
  • Sinusoidal schedule
    $1\displaystyle 2r(n)=\sin{\left[(2\times ratio-1)\times\frac{\pi}{2}\right]}+$ 1 (6)
  • Linear schedule
    $r(n)=ratio$ (7)
E.2.3 Ratio variants
  • Logarithmic ratio
    $\frac{\log{n}}{\log{N}}$ (8)
  • Linear ratio
    $\frac{n}{N}$ (9)
E.3 Ablation study results
We conduct experiments for all 8 variants on DoorKey and CleanHouse for 8 seeds. The results shown in Figure 14 present no significant differences between all 8 variants.
(a) DoorKey
(b) CleanHouse
Figure 14: Comparing Scheduled HC variants. We compare our proposed Scheduled HC with its variants. The label stands for interpolation-schedule-ratio. For example, when using logarithmic interpolation, sinusoidal schedule, and logarithmic ratios, our label is Log-Sin-Log. The results indicate no significant differences between all Scheduled HC variants.
@@ Appendix F Baselines and their hyperparameters
We compare our framework to four baselines, LEAPS [Trivedi et al., 2021], HPRL [Liu et al., 2023], CEBS [Carvalho et al., 2024], and HC [Carvalho et al., 2024]. LEAPS, HPRL, and CEBS search programs in the latent spaces, while HC searches programs in the programmatic space. We first introduce the concept of latent space in Section F.1, then dive into the details of each baseline in the following sections. The pseudo-code of CEBS is described in Algorithm 1, and the pseudo-code of HC is described in Algorithm 2.
F.1 Latent space
In addition to the programmatic space, another choice is to search in a learned latent space [Trivedi et al., 2021, Liu et al., 2023], constructed by training a variational autoencoder on randomly generated DSL programs. To sample a neighbor program, a small noise is added to the latent embedding of the current program candidate, and the decoder can generate a program from the noise-corrupted embedding.
F.2 LEAPS
LEAPS utilizes the Cross-entropy method (CEM) [Rubinstein and Kroese, 2004] as its search strategy. It first generates $k$ neighborhood programs from its latent space and evaluates their episodic return on the Karel environment [Trivedi et al., 2021]. Next, if the average return over the top-$E$ neighborhood programs is higher than the best-seen average return, it continues the process on the mean latent embedding of the top-$E$ neighborhood programs. When the average return no longer increases, the program with the highest episodic return is returned.
@@ F.2 LEAPS
F.3 Hierarchical programmatic reinforcement learning (HPRL)
@@
F.4 Cross-entropy beam search (CEBS)
CEBS [Carvalho et al., 2024] extends CEM to maintain a set of $E$ candidate programs to perform beam search. Unlike CEM which only maintains one candidate program. CEBS maintains a set of $E$ candidate programs, to perform beam search. In other words, CEBS searches all the neighborhoods of the top-$E$ programs.
It also utilizes the pre-trained VAE weight from LEAPS [Trivedi et al., 2021] to search the program in the latent space. All the hyperparameter follows the original CEBS with neighborhood $K$ equal to 64, candidate $E$ equal to 16, and noise ratio $\sigma$ equal to 0.25 for all ten Karel tasks. The pseudo-code is described in Algorithm 1.
F.5 Hill climbing (HC)
@@ Appendix H Aggregated performance across all tasks
Appendix I Wall-clock time evaluation on task DoorKey
We provide a wall-clock time evaluation (real-time elapsed) for the DoorKey task in Figure 18, with the x-axis showing the number of seconds on a log scale. It also accounts for the time LLM takes to generate the programs. The result highlights the superiority of our proposed LLM-GS framework.
Figure 18: Wall-clock time evaluation on the DoorKey task. We evaluate the wall-clock time performance on the DoorKey task, with the x-axis representing the number of seconds on a logarithmic scale. The plot also includes the time LLM takes to generate the programs. The results demonstrate the superiority of our proposed LLM-GS framework.
(a) StairClimber
(b) Maze
(c) FourCorners
(d) TopOff
(e) Harvester
(f) CleanHouse
(g) DoorKey
(h) OneStroke
(i) Seeder
(j) Snake
Figure 19: Statistical test on the Karel and Karel-Hard tasks. We use the Student’s t-test to measure the mean of LLM-GS and HC. The p-value of the t-test (red line) follows the left axis, while LLM-GS and HC use the right axis. The brown dotted line presents the significant level of the null hypothesis. LLM-GS outperforms HC on 8 / 10 tasks when the search ends.
Appendix J Statistical test between LLM-GS and HC
Since the confidence interval of LLM-GS overlaps with that of HC in some tasks, we conduct hypothesis testing between these two methods for better comparison. The null hypothesis $H_{0}$ is that the mean of LLM-GS is greater than HC. We use Student’s t-test [STUDENT, 1908] to test the distribution of LLM-GS and HC at the same number of evaluated programs.  Figure 19 shows the p-value from the t-test across the evaluated program on the Karel and Karel-hard tasks. We adjust the primary y-axis (left) and the secondary y-axis (right) to have the same scale. The brown dotted line is 0.05. If the red lines (p-value) are under the brown dotted line, LLM-GS statistically outperforms HC. If the red lines are above the brown dotted line, LLM-GS fails to surpass HC.
The result shows that our method significantly outperforms HC on StairClimber, Maze, FourCorners, DoorKey, and Seeder. On task TopOff, LLM initialization can not outperform HC in the first few attempts but has superior efficiency when the search ends. On the task Harvester, LLM-GS can surpass HC in less than a thousand programs, but cannot find a difference at the end of the search. On the task CleanHouse, LLM-GS cannot outperform HC except for the LLM initialization. On the task OneStroke and Snake, LLM-GS eventually surpasses HC when the search process ends. The LLM initialization (32 is at the middle point of 10 and 100) surpasses HC on all tasks. On the other hand, LLM-GS outperforms HC on 8 / 10 tasks (except for Harvester and CleanHouse) when the search ends.
Appendix K LLM prompts for ablations and revision
@@ Appendix I Wall-clock time evaluation on task DoorKey
Ablation prompts for program generation methods. We conduct an ablation study in Section 5.3 to justify that our LLM-generating program method performs best in acceptance rate and best return. Here we list the complete prompts. There are three approaches to generating DSL programs: Pythonic-DSL, Python, and DSL. There are two types of prompt, system prompt and user prompt, in all approaches. In the system prompt, both the Python and the Pythonic-DSL approaches contain limitations of Python usage. The DSL approach contains the grammar of the Karel, and the Pythonic-DSL contains the paired Python-like and Karel production rules. All of the system prompts contain the environment physics, action, and perception. On the other hand, the user prompt contains five placeholders for task name, map description, initial position, task goal, and task reward. In Appendix C, all of the task-dependent information can be filled in the placeholders. Section D.2 shows the full Pythonic-DSL prompt with the task DoorKey filled in the placeholders. The Pythonic-DSL prompts are in Section K.1, and the Python prompts are in Section K.2, and the DSL prompts are in Section K.3.
@@
Prompts for LLM revision. We list all of the user prompts in the experiment of LLM revision, the system prompt is the same as the one in Section K.1. We implement four approaches to revising the program: Regenerate, Regenerate with reward, Agent execution trace, and Agent and program execution trace. We ask the LLM to regenerate with all generated programs in the last round without repetition, and the Regenerate prompt is in Section K.4. We ask the LLM to regenerate with the program relating to reward, and the revision prompt of the Regenerate with reward is in Section K.5. We utilize the execution trace of the Karel agent in the grid world, and the revision prompt of the Agent execution trace is in Section K.6. We provided the action/perception call and executing line, and the revision prompt of the Agent and program execution trace is in Section K.7. The program Figure 20 is the revision target of our method revision method in Section K.6 and Section K.7. Figure 20 is the trajectories of the program we used in method Section K.6 and Section K.7. In the example, the origin program has an average return of 0.5. LLM revision with agent trace can reach 0.640625, while LLM revision with both agent and program traces can reach a result of 0.8125.
@@
[!ht] {minted}Python def run(): while noMarkersPresent(): if frontIsClear(): move() else: turnLeft() pickMarker() while frontIsClear(): move() turnRight() while noMarkersPresent(): if frontIsClear(): move() else: turnLeft() if frontIsClear(): move() turnRight() putMarker() The revision results of Section K.6.
@@
[!ht] {minted}Python def run(): while noMarkersPresent(): if frontIsClear(): move() else: turnLeft() pickMarker() while noMarkersPresent(): if rightIsClear(): turnRight() move() else: if frontIsClear(): move() else: turnLeft() putMarker() The revision results of Section K.7.
K.1 Pythonic-DSL
K.1.1 System prompt
@@
K.2 Python
K.2.1 System prompt
@@
K.5 Regenerate with reward
@@
K.6 Agent execution trace
@@
K.7 Agent and Program execution trace.
@@
Appendix L Minigrid Environment
Minigrid [Chevalier-Boisvert et al., 2023] is a grid-world environment designed for conducting research on reinforcement learning [Guan et al., 2022, Jelley et al., 2023, Furelos-Blanco et al., 2023]. Unlike the Karel environment [Pattis, 1981], Minigrid has a richer set of objects, and each object comes with a color. Besides Karel’s actions, the Minigrid agent also can ”toggle” objects, like unlocking a door or opening a box. The environment has built-in tasks that range from simple navigation to complex multi-step reasoning.
To evaluate our proposed LLM-GS framework on the Minigrid environment, we designed and implemented a domain-specific language (DSL), which contains more actions and object types than the Karel one. We will introduce the DSL in Section L.1, the tasks in Section L.2, the result in Section L.3, and the prompts in Section L.4.
L.1 MiniGrid DSL
Here, we present the DSL we designed for the Minigrid environment. The actions are similar to the ones from Karel. The perceptions, on the other hand, are distinct from Karel. Specifically, the perception functions in our Minigrid DSL are parameterized, i.e., require an input parameter such as an object, since there are multiple object types in the domain, unlike the case in Karel. Thus, we implemented two novel perceptions: front_object_type( type ) and front_object_color( color ), which are completely different from the ones in Karel DSL. They check if the cell right in front of the agent has a certain type or a certain color object. We list the Python to DSL converting rules in Table 9 and the probabilities of the production rules in Table 10. The setting for the program mutation is the same as the one in Karel DSL.
Table 9: Python to DSL converting rules in Minigrid.
Python DSL
def run(): s DEF run m( s m)
while b: s WHILE c( b c) w( s w)
if b: s IF c( b c) i( s i)
if b: s else: s IFELSE c( b c) i( s i) ELSE e( s e)
for i in range(n): s REPEAT R=n r( s r)
not h not c( h c)
front_is_clear() front_is_clear
front_object_type(type) front_object_type h( type h)
front_object_color(color) front_object_color h( color h)
is_carrying_object() is_carrying_object
forward() forward
left() left
right() right
pickup() pickup
drop() drop
toggle() toggle
Table 10: Probabilities of the production rules in Minigrid.
Category Rule Probability
Program P Statement 1.0
Statement S While 0.15
Repeat 0.03
Concatenate 0.5
If 0.08
Ifelse 0.04
Action 0.2
Condition c Boolean 0.5
not 0.5
Action a forward 0.5
left 0.15
right 0.15
pickup 0.08
drop 0.08
toggle 0.04
Boolean b front_is_clear 0.125
front_object_type 0.5
front_object_color 0.25
is_carrying_object 0.125
Type t lava 0.25
door 0.25
ball 0.25
box 0.25
Color o blue 0.5
red 0.5
Number n i (0 $\leq$ i $\leq$ 19) 0.05
L.2 Minigrid tasks
We evaluate three tasks from the Minigrid environment: Lava Gap (Section L.2.1), Put Near (Section L.2.2), and Red Blue Door (Section L.2.3). We include an example illustration of the three tasks in Figure 21 and introduce the details of each task in the following sections.
(a) LavaGap
(b) PutNear
(c) RedBlueDoor
Figure 21: Illustrations of the initial and final state of the task we chose for the Minigrid [Chevalier-Boisvert et al., 2023] evaluation. The position of obstacles, objects, and the agent’s position are randomly set according to the configurations of each task.
L.2.1 LavaGap
In the task LavaGap, the agent must pass through a gap in a vertical strip of lava and reach the goal. Stepping on the lava will terminate the episode. We set the reward function of LavaGap to give a reward of -1.0 when stepping on lava and a reward of 1.0 when reaching the goal.
Purpose Prompt Text
Task Name LAVAGAP
Map Description The map is a 6x6 grid surrounded by walls. There is a vertical strip of lava with a randomly selected gap in the middle of the map and a goal square at the bottom right corner.
Initial Position The agent starts at the top left corner of the map facing east.
Task Goal The agent has to reach the goal square at the bottom right corner of the map without step on lava. Touching the lava terminates the episode.
Task Reward If the agent reaches the goal, it will receive a reward of 1. If the agent steps on lava, it will receive a reward of -1.
L.2.2 PutNear
In the task PutNear, the agent needs to pick up the move object and drop it next to the target object. We modified the task to fix the move object as a ball and the target object as a box. Picking up the move object will get a reward of 0.5 and dropping it next to the target object will receive the remaining 0.5. Picking up the wrong object will get a reward of -1.0.
Purpose Prompt Text
Task Name PUTNEAR
Map Description The map is a 6x6 grid surrounded by walls. There are two objects, one is a ball and the other is a box. The two objects are randomly placed on the map.
Initial Position The agent starts at the top left corner of the map facing east.
Task Goal The agent has to first locate and pick up the ball. Next, it needs to find the box and drop the ball either to the right or left of it. Picking up the wrong object will terminate the episode.
Task Reward If the agent picks up the ball, it will receive a reward of 0.5. If the agent drops the ball to the right or the left of the box, it will receive a reward of 1. If the agent picks up the wrong object, it will receive a reward of -1.
L.2.3 RedBlueDoor
In the task RedBlueDoor, the agent must first open the red door and then the blue one. Opening the red door first will get a reward of 0.5, followed by an additional 0.5 for opening the blue door afterward. Opening the blue door before the red one will get a reward of -1.0.
Purpose Prompt Text
Task Name REDBLUEDOOR
Map Description The map is a 6x12 grid surrounded by walls that are vertically split into three chambers. The left chamber is a 4x2 grid, the middle chamber is a 4x4 grid and the right chamber is a 4x2 grid. On the wall between the left and middle chamber, there is a randomly placed red door. On the wall between the middle and right chamber, there is a randomly placed blue door.
Initial Position The agent starts at a random position with a random direction in the middle chamber.
Task Goal The agent has to first find and open the red door, then find and open the blue door. Opening the blue door first will terminate the episode.
Task Reward If the agent opens the red door first, it will receive a reward of 0.5. If the agent opens the blue door after opening the red door, it will receive a reward of 0.5. If the agent opens the blue door first, it will receive a reward of -1.
L.3 Minigrid domain
L.3.1 Evaluation setup
We evaluate our proposed LLM-GS framework using the Minigrid tasks LavaGap, PutNear, and RedBlueDoor. More task details can be found in Section L.2. LLM-GS is compared with the best-performing baseline HC. The evaluation metric is the same as the one mentioned in Section 5.1 except for $C=16$ and $N=10^{4}$. For the setup, we use GPT-4o (gpt-4o-2024-08-06 with temperature=1.0, top_p=1.0) as our LLM module to generate the initial search population. The scheduler of our proposed Scheduled HC starts from $K_{start}=32$ to $K_{end}=2048$. We evaluate our LLM-GS and HC with 8 random seeds.
L.3.2 Experimental results
Figure 22 presents the experimental result, indicating that our proposed LLM-GS is more sample-efficient compared to HC on all the tasks.
(a) LavaGap
(b) PutNear
(c) RedBlueDoor
Figure 22: Efficiency in Minigrid tasks. We compare our proposed LLM-GS framework against best-performing baseline HC in three Minigrid tasks LavaGap, PutNear, and RedBlueDoor. The results show that our LLM-GS is more sample-efficient than HC.
L.4 Minigrid prompts
L.4.1 The system prompt
In the Minigrid environment, you navigate a grid-based world referred to as a "map." Within this map, an entity called the "agent" can move, change direction, and interact with objects. The map also includes obstacles like "walls," which block the agents movement; if the agent encounters a wall, it remains stationary. Another obstacle is "lava," which ends the episode if the agent steps on it. Additionally, there are "doors" that block the agent until opened using the "toggle" action.
At the start of each episode, various objects are scattered across the map, and the agent can interact with them as needed. These objects fall into four categories: "lava," "door," "ball," and "box." Except for lava, objects can appear in two colors: "red" and "blue."
Lava: Ends the episode when stepped on.
Door: Blocks the agents path until opened using the "toggle" action.
Ball: Can be picked up and dropped.
Box: Can be picked up and dropped.
Here are the available actions for the agent:
forward(): Asks the agent to move forward one cell. The agent will stay still if a wall or an unopened door is blocking its way. The episode will be terminated if the agent steps on lava.
left(): Asks the agent to rotate 90 degrees counterclockwise.
right(): Asks the agent to rotate 90 degrees clockwise.
pickup(): Asks the agent to pick up an object right in front of the current cell. The agent can only carry one object at a time.
drop(): Asks the agent to drop the object it is carrying right in front of the current cell. The agent can only drop the object if it is carrying one and there is no object in the cell.
toggle(): Asks the agent to change the state of a door directly in front of its current position. The agent can use this action to open or close the door immediately ahead.
Here are the available perceptions of the agent:
front_is_clear(): Returns True if there is no wall and an unopened door right in front of the agent. It can only check wall and unopened door, so if there is lava right in front of the agent, it will return True.
front_object_type(type: Literal["lava", "door", "ball", "box"]): Returns True if there is an object of the specified type right in front of the agent. The type can be and only can be "lava", "door", "ball", or "box".
front_object_color(color: Literal["red", "blue"]): Returns True if there is an object of the specified color right in front of the agent. The color can be and only can be "red" or "blue".
is_carrying_object(): Returns True if the agent is carrying an object.
There are some limitations to the Python program:
- do not define other functions besides run()
- do not call other functions
- do not define variables
- do not use True, False, break, continue, return, ==, !=, elif, or, and
Python to Minigrid dsl conversion
1. "def run(): s" to "DEF run m( s m)"
2. "while b: s" to "WHILE c( b c) w( s w)"
3. "if b: s" to "IF c( b c) i( s i)"
4. "if b: s else: s" to "IFELSE c( b c) i( s i) ELSE e( s e)"
5. "for i in range(n): s" to "REPEAT R=n r( s r)"
6. "not h" to "not c( h c)"
7. "front_is_clear()" to "front_is_clear"
8. "front_object_type(type)" to "front_object_type h( type h)"
9. "front_object_color(color)" to "front_object_color h( color h)"
10. "is_carrying_object()" to "is_carrying_object"
11. "forward()" to "forward"
12. "left()" to "left"
13. "right()" to "right"
14. "pickup()" to "pickup"
15. "drop()" to "drop"
16. "toggle()" to "toggle"
L.4.2 The user prompt of the task LavaGap
Ill provide you with the task name and description.
Task name: LAVAGAP
Task map: The map is a 6x6 grid surrounded by walls. There is a vertical strip of lava with a randomly selected gap in the middle of the map and a goal square at the bottom right corner.
Task agent position: The agent starts at the top left corner of the map facing east.
Task goal: The agent has to reach the goal square at the bottom right corner of the map without step on lava. Touching the lava terminates the episode.
Task return: If the agent reaches the goal, it will receive a reward of 1. If the agent steps on lava, it will receive a reward of -1.
2. Convert the Python program to the Minigrid dsl program.
L.4.3 The sample response of LLM in the task LavaGap
‘‘‘python
def run():
while not front_object_type("lava"):
forward()
left()
forward()
right()
while front_is_clear():
forward()
right()
forward()
left()
while front_is_clear():
forward()
Minigrid DSL program:
DEF run m(
WHILE c( not c( front_object_type h( "lava" h) c) c) w(
forward
w)
left
forward
right
WHILE c( front_is_clear c) w(
forward
w)
right
forward
left
WHILE c( front_is_clear c) w(
forward
w)
m)
Appendix M Limitation